OpenStudy (samigupta8):

Two parallel plate capacitor with diff distance between the plates are connected in parallel to a voltage source. A point positive charge is moved from a point 1 that is exactly in middle between the plates of capacitor C1 to a point 2 in capacitor C2 that lies at a distance from the negative plate of C2 equal to half the distance between the plates of C1.Then the work done by electric force is

OpenStudy (samigupta8):

@ganeshie8

OpenStudy (samigupta8):

Options are A).positive B).negative C).0 D).data sufficient

OpenStudy (samigupta8):

Oops insufficient

ganeshie8 (ganeshie8):

any picture given ?

OpenStudy (samigupta8):

Yep !

OpenStudy (samigupta8):

Bt i don't have option to draw...

ganeshie8 (ganeshie8):

you may attach the file..

OpenStudy (samigupta8):

No such option here .. Only post is written...

ganeshie8 (ganeshie8):

you're on mobile ?

OpenStudy (samigupta8):

Yep..

ganeshie8 (ganeshie8):

tell me the name of your textbook

OpenStudy (samigupta8):

Actually the diagram is just depicting the situation given in the question and nothing else..

OpenStudy (samigupta8):

R u able to draw the diagram requested in the question???

OpenStudy (michele_laino):

the situation of the problem can be represented by this drawing:  so the work done by the electric field, is: \[{L_E} = \left( {4\pi {\sigma _1}\frac{{{d_1}}}{2} - 4\pi {\sigma _2}\frac{{{d_1}}}{2}} \right)q = 2\pi {d_1}\left( {{\sigma _1} - {\sigma _2}} \right)q\] where \(q>0\), \(d_1,\;d_2\) are the separation of the capacitor \(C_1,\;C_2\) respectively \(V\) is the external voltage Next, we can write: \[{\sigma _1} - {\sigma _2} = \frac{{\left( {{\sigma _1} - {\sigma _2}} \right)S}}{S} = \frac{{\left( {{Q_1} - {Q_2}} \right)}}{S} = \frac{{\left( {{C_1} - {C_2}} \right)V}}{S}\] furthermore, we have: \[2\pi {\sigma _1} = \frac{S}{{2{C_1}}}\] so, after a substitution, we get: \[\begin{gathered} {L_E} = \left( {4\pi {\sigma _1}\frac{{{d_1}}}{2} - 4\pi {\sigma _2}\frac{{{d_1}}}{2}} \right)q = 2\pi {d_1}\left( {{\sigma _1} - {\sigma _2}} \right)q \hfill \\ \hfill \\ = \frac{S}{{2{C_1}}} \cdot \frac{{\left( {{C_1} - {C_2}} \right)V}}{S} \cdot q = \boxed{\frac{{\left( {{C_1} - {C_2}} \right)qV}}{{2{C_1}}}} \hfill \\ \end{gathered} \]

OpenStudy (samigupta8):

How do we have this ? S/2C1=2πsigma1

OpenStudy (samigupta8):

And sir you considered just the potential at both points due to one plate of cPacitor and not due to other.. Why so??